If the length of the hypotenuse of a rigth triangle is 15 inches, and the length of one side is 9 inches, what is the length of the other side?
Answer by fractalier(6550)

9^2 + x^2 = 15^2
81 + x^2 = 225
x^2 = 144
x = 12